# Experimental single-copy distillation of quantumness from higher-dimensional entanglement

### v1

@@ Abstract
is larger than $F^{c}_{d}=\frac{1}{d}$, where $|\Psi_{d}\rangle=(\mathbb{I}_{d}\otimes U_{d})|\Phi^{+}_{d}\rangle$ is an arbitrary maximally entangled two-qudit state, $|\Phi^{+}_{d}\rangle=\frac{1}{\sqrt{d}}\sum_{i=0}^{d-1}|i\rangle|i\rangle$, while $\mathbb{I}_{d}$ and $U_{d}$ are the identity and an arbitrary unitary matrix acting on $\mathbb{C}^{d}$, respectively. From Eq. 3, one can easily verify that if a two-qubit state $\rho$ satisfies $F_{2}>\frac{1}{2}$, it must also satisfy $F_{d}>\frac{1}{d}$ for all integer $d>2$. Since all two-qubit states violating the CHSH-Bell inequality satisfy Horodecki et al. (1996) $F_{2}>\frac{1}{2}$, the CHSH-Bell violation of $\rho_{\text{expt},f}(v)$ for $v\leq 0.15$ already confirm their usefulness for teleporting a qubit state, and hence an arbitrary qudit state. In fact, by using Eq. 3 and the explicit form of $\rho_{\text{expt},f}(v)$, we see from Fig. 3b that the teleportation power of all $\rho_{\text{expt}}(v)$ with $v<v_{\text{distill}}$ is successfully recovered via our qubit ScLF, in agreement with the results of Li et al. (2021). In Section B.3, we show that the same ScLF also activates the dense-codability Bruß et al. (2004) of some $W^{(d)}(v)$ and provide experimental evidence in Section C.3 confirming these findings.
Finally, let us compare the steering robustness Uola et al. (2020) SR—a quantifier of steerability monotonously related Piani and Watrous (2015) to the usefulness of a state for some specific subchannel discrimination task—between $\rho_{\text{expt}}(v)$ and $\rho_{\text{expt},f}(v)$. To our knowledge, the exact value of $\text{SR}[W^{(d)}(v)]$ is unknown. In Appendix B.2, we provide the best lower bound on $\text{SR}[W^{(d)}(v)]$ that we have found using the technique of Cavalcanti and Skrzypczyk (2017) by considering up to $n_{s}=7$ measurement bases. Accordingly, we see in Fig. 4 an improvement from these lower bounds on $\text{SR}[\rho_{\text{expt}}(v)]$ to $\text{SR}[\rho_{\text{expt},f}(v)]$ for all $v<0.25$. In particular, the steerability of $\rho_{\text{expt},f}(v)$ for $v=0.2$ is clearly witnessed while that of $\rho_{\text{expt}}(v)$ is unknown. Note further from Fig. 4 that, in principle, steerability may also be activated beyond $v>v_{\text{steer}}=\frac{2}{9}$ for $d=3$. However, due to experimental imperfections, we do not observe the activation for $v=0.25$.
@@
Conclusion.–Using qubit local filtering operations, we experimentally demonstrate the single-copy distillation of several genuine quantum features from the paradigmatic family of higher-dimensional (specifically two-qutrit) Werner states. Notice that via the coordinated local operations of $U\otimes U$-twirling, we can always transform any bipartite quantum state to a Werner state. Thus, in principle, our experimental demonstration of single-copy distillation also applies to all quantum states that transform to a 1-distillable two-qutrit Werner state by these local operations. In turn, our work also demonstrates the single-copy operation involved in the entanglement distillation of higher-dimensional quantum states via Werner states.
@@
To facilitate the experimental preparation of Werner states, we provide an intuitive decomposition of any finite-dimensional Werner states as a mixture of two-qubit states (see also Section A.5). The decomposition, in turn, sheds light on the relevance of qubit filtering operations for these states, partially addressing one of the open problems from Li et al. (2021). Conceivably, our scheme for converting qubit states to higher-dimensional states may also find applications in the experimental preparation of other higher-dimensional states admitting a similar two-qubit decomposition.
@@
Among the various quantum features distilled from the experimentally prepared two-qutrit Werner states, their Bell nonlocality is especially noteworthy. Indeed, our work provides the first proof-of-principle certification of the Bell nonlocality of these important higher-dimensional entangled states, hidden or otherwise. Despite experimental imperfections, we have also recovered the teleportation power for the experimentally prepared 1-distillable two-qutrit Werner states. Using the lower bound on steering robustness we have established theoretically (see Section B.2), our experimental results also provide strong evidence for enhancing steering robustness (and hence usefulness for certain discrimination tasks Piani and Watrous (2015)) via single-copy distillation. Note that the former theoretical results may also be of independent interest, e.g., in the studies of measurement incompatibility Chen et al. (2016).
For future work, within the realm of single-copy local filtering operations, it is clearly of interest to experimentally demonstrate the hidden nonlocality of the five-dimensional Werner states, first discussed in Popescu (1995), and the hidden steerability of the three-dimensional Werner states we barely miss in the current experiment. Since the filtered two-qubit state only exhibits a tiny amount of Bell-nonlocality (steerability), the challenges here are to produce very high-quality, high-dimensional Werner states and follow up with extremely well-controlled qubit projections (and measurements). Another closely related problem worth exploring is the power of local filtering operations beyond a qubit projection.
For example, could a higher-dimensional local filtering operation be more effective, e.g., in demonstrating the hidden nonlocality of $W^{(d)}(v)$ for a broader range of the parameter $v$? A systematic investigation of the problem is surely welcome. Similarly, suppose a more efficient high-dimensional local filtering operation exists to activate dense codability. In that case, the advantage gained in these operations might outweigh the need for classical communication in implementing such operations, thereby overcoming the objection raised in Bruß et al. (2004) for allowing such operations in the characterization of dense-codability. An analytic proof confirming our numerical observation of when a Werner state admits a quasi-symmetric extension or a bosonic symmetric extension is desirable, too.
@@ Acknowledgements.
We thank Andrew Doherty, Huan-Yu Ku, and Paul Skrzypczyk for useful discussions. KSC thanks the hospitality of the Institut Néel, where part of this work was completed. This work is supported by the National Key R&D Program of China (Grants No. 2019YFA0308200), the Shandong Provincial Natural Science Foundation (Grant No. ZR2023LLZ005), the Taishan Scholar of Shandong Province (Grants No. tsqn202103013), the Shenzhen Fundamental Research Program (Grant No. JCYJ20220530141013029), the National Science and Technology Council, Taiwan (Grants No. 109-2112-M-006-010-MY3, 112-2628-M-006-007-MY4, 113-2917-I-006-023, 113-2918-I-006-001) and the Foxconn Research Institute, Taipei, Taiwan. This research was supported in part by the Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported by the Government of Canada through the Department of Innovation, Science, and Economic Development, and by the Province of Ontario through the Ministry of Colleges and Universities.
@@ A.1 Entanglement source
Figure 5: Zoom-in view of the “Entanglement source” part of the experimental setup shown in Fig. 2a for generating maximally entangled photon pairs.
@@
For our experiment, the power intensity of pump light is set at $5$ mW, and we observe a twofold coincidence count rate of 73 kHz. The two photons are coupled out by two collimators, and the polarization of each photon is corrected by polarization controller (PC) consisting two QWPs and a HWP. Finally, we let photon B go through an additional HWP at 90 to produce the singlet state $\varrho_{0}^{H,V}=(|\Psi^{-}\rangle\!\langle\Psi^{-}|)_{\text{AB}}$, where $|\Psi^{-}\rangle_{\text{AB}}=\frac{1}{\sqrt{2}}|H\rangle|V\rangle-|V\rangle|H\rangle)_{\text{AB}}$.
@@ A.5 Experimental realization of an arbitrary Werner state
Although the decomposition that we have provided in Eq. 2 suffices for our experimental demonstration, it is not valid for $v>\frac{d+1}{2d}$, which leads to a negative value of $q=1-\frac{2d}{d+1}v$. For completeness, we provide below an alternative two-qudit decomposition of $W^{(d)}(v)$ valid for all $v\in[0,1]$. Explicitly, we note that
@@ Appendix B Miscellaneous Details and Theoretical Results
B.1 Bell-nonlocality
B.1.1 Symmetric (quasi-)extension
@@
For completeness (and simplicity of presentation), we provide below a semidefinite program (SDP) that facilitates the search for the existence of a $(k,1)$-symmetric extension Doherty et al. (2002) for any given $\rho_{\text{AB}}$ acting on $\mathbb{C}^{d_{A}}\otimes\mathbb{C}^{d_{B}}$. For the rest of this subsection, we shall refer to any $\rho_{\text{AB}}$ admitting a $(k,1)$-symmetric extension as being $k$-extendible. Our SDP formulation differs somewhat from of Doherty et al. (2002); Terhal et al. (2003) but makes explicit use of the two facts: (1) the set of $k$-extendible density matrices $\rho_{\text{AB}}$ is convex, and (2) the maximally mixed state $\frac{\mathbb{I}_{D}}{D}$, where $D:=d_{A}d_{B}$ is $k$-extendible for all $k>1$. Hence, if mixing $\rho_{\text{AB}}$ with a negatively-weighted $\frac{\mathbb{I}_{D}}{D}$ is $k$-extendible, $\rho_{\text{AB}}$ itself must already be $k$-extendible:
@@
For Werner states $W^{(d)}(w)$, using the identity $\Pi^{d}_{+}=\mathbb{I}-\Pi^{d}_{-}$, we can also rewrite it from Eq. 1 as
$v_{t}=\frac{n^{d}_{+}}{D}w_{t}=\frac{n^{d}_{+}}{D}\frac{t^{\star}_{0}-1}{t^{\star}_{0}},$ (18)
i.e., $W^{(d)}(v)$ for all $v>v_{t}$ is $k$-extendible.
$\ \ \min_{t\geq 0}\qquad\qquad\qquad\qquad t$
@@
B.1.2 Finding Bell-nonlocal correlations for the experimentally produced two-qutrit states

### v2

@@ Abstract
is larger than $F^{c}_{d}=\frac{1}{d}$, where $|\Psi_{d}\rangle=(\mathbb{I}_{d}\otimes U_{d})|\Phi^{+}_{d}\rangle$ is an arbitrary maximally entangled two-qudit state, $|\Phi^{+}_{d}\rangle=\frac{1}{\sqrt{d}}\sum_{i=0}^{d-1}|i\rangle|i\rangle$, while $\mathbb{I}_{d}$ and $U_{d}$ are the identity and an arbitrary unitary matrix acting on $\mathbb{C}^{d}$, respectively. From Eq. 3, one can easily verify that if a two-qubit state $\rho$ satisfies $F_{2}>\frac{1}{2}$, it must also satisfy $F_{d}>\frac{1}{d}$ for all integer $d>2$ (see Section B.1). Since all two-qubit states violating the CHSH-Bell inequality satisfy Horodecki et al. (1996) $F_{2}>\frac{1}{2}$, the CHSH-Bell violation of $\rho_{\text{expt},f}(v)$ for $v\leq 0.05$ already confirms, modulo statistical uncertainty and the independent and identical trials assumption, their usefulness for teleportation. In fact, by using Eq. 3 and the explicit form of $\rho_{\text{expt},f}(v)$, we see from Fig. 3b that the teleportation power of all $\rho_{\text{expt}}(v)$ with $v<v_{\text{distill}}$ is successfully recovered via our qubit ScLF, in agreement with the results of Li et al. (2021). In Section B.4, we show that the same ScLF also activates the dense-codability Bruß et al. (2004) of some $W^{(d)}(v)$ and provide experimental evidence in Section C.3 confirming these findings.
Finally, we compare the steering robustness Uola et al. (2020) SR—a quantifier of steerability monotonously related Piani and Watrous (2015) to the usefulness of a state for some channel discrimination task—between $\rho_{\text{expt}}(v)$ and $\rho_{\text{expt},f}(v)$. To our knowledge, the exact value of $\text{SR}[W^{(d)}(v)]$ is unknown. In Section B.3, we provide the best lower bound on $\text{SR}[W^{(d)}(v)]$ we have found Cavalcanti and Skrzypczyk (2017) by considering up to $n_{s}=7$ measurement bases. Accordingly, we see in Fig. 4 an improvement from these lower bounds on $\text{SR}[\rho_{\text{expt}}(v)]$ to $\text{SR}[\rho_{\text{expt},f}(v)]$ for all $v<0.25$. In particular, the steerability of $\rho_{\text{expt},f}(v)$ for $v=0.2$ is clearly witnessed while that of $\rho_{\text{expt}}(v)$ is unknown. Note further from Fig. 4 that, in principle, steerability may also be activated beyond $v>v_{\text{steer}}=\frac{2}{9}$ for $d=3$. However, due to experimental imperfections, we do not observe the activation for $v=0.25$.
@@
Conclusion.–Using local qubit filtering operations, we experimentally demonstrate the single-copy distillation of several quantum features from the paradigmatic family of higher-dimensional (two-qutrit) Werner states. Via the coordinated local operations of $U\otimes U$-twirling, one can always transform any bipartite quantum state to a Werner state. Thus, our experimental demonstration of single-copy distillation also applies to all quantum states that transform to a 1-distillable two-qutrit Werner state by these local operations. In turn, this also demonstrates a key step involved in the conventional entanglement distillation Horodecki and Horodecki (1999) from higher-dimensional quantum states via Werner states.
@@
To experimentally prepare Werner states, we provide an intuitive decomposition of any finite-dimensional Werner states as a mixture of two-qubit states (see also Section A.5). The decomposition, in turn, sheds light on the relevance of qubit filtering operations for these states, partially addressing one of the open problems from Li et al. (2021). Conceivably, our experimental scheme for converting qubit states to higher-dimensional states may also find applications in preparing other higher-dimensional states.
@@
Among the various quantum features distilled from the experimentally prepared two-qutrit Werner states, their Bell nonlocality is especially noteworthy. Indeed, our work provides the first proof-of-principle validation of the Bell nonlocality of these important higher-dimensional states, hidden or otherwise. Despite experimental imperfections, we have also recovered the missing teleportation power in some of these higher-dimensional states. Using the lower bound on steering robustness we have established theoretically (see Section B.3), our experimental results also provide strong evidence for steering robustness enhancement (and hence usefulness for certain discrimination tasks Piani and Watrous (2015)) via single-copy distillation. The former theoretical results may also be of independent interest, e.g., in the studies of measurement incompatibility Chen et al. (2016).
Note again that, in contrast with other entanglement distillation protocols Bennett et al. (1996a); Deutsch et al. (1996); Bennett et al. (1996b); Horodecki and Horodecki (1999), our ScLFs only require local manipulations at individual copies of the noisy states, thus making them practical, especially in a photonic setup. Indeed, given its simplicity and effectiveness, it is surely worth exploring their implementation also in other quantum information processing platforms, e.g., solid state systems. For future work, exploring the power of ScLFs beyond a qubit projection is clearly of interest. For example, could a higher-dimensional local filtering operation be more effective, e.g., in demonstrating the hidden nonlocality of more $W^{(d)}(v)$ or other states? If they exist, we can always experimentally implement them (see  Section A.6), transforming a $d$-dimensional state to a $d^{\prime}$-dimensional state (for any $d^{\prime}\leq d$). Similarly, if such operations for activating dense codability exist, the advantage gained in implementing them might outweigh the need for classical communication involved in their realization, thereby overcoming the objection raised in Bruß et al. (2004) for allowing such operations in the characterization of dense-codability. However, since Werner states admit the qubit decomposition of Eq. 2, we conjecture that a more effective HD filter does not exist. An analytic proof confirming our numerical observation of when a Werner state admits a quasi-symmetric extension or a bosonic symmetric extension is desirable, too.
@@ Acknowledgements.
Acknowledgments.– We thank Andrew Doherty, Huan-Yu Ku, Paul Skrzypczyk, and Yujie Zhang for useful discussions. KSC thanks the hospitality of the Institut Néel, where part of this work was completed. This work is supported by the National Key R&D Program of China (Grants No. 2019YFA0308200), the Shandong Provincial Natural Science Foundation (Grant No. ZR2023LLZ005), the Taishan Scholar of Shandong Province (Grants No. tsqn202103013), the Shenzhen Fundamental Research Program (Grant No. JCYJ20220530141013029), the National Science and Technology Council, Taiwan (Grants No. 109-2112-M-006-010-MY3, 112-2628-M-006-007-MY4, 113-2917-I-006-023, 113-2918-I-006-001) and the Foxconn Research Institute, Taipei, Taiwan. This research was supported in part by the Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported by the Government of Canada through the Department of Innovation, Science, and Economic Development, and by the Province of Ontario through the Ministry of Colleges and Universities.
@@ A.1 Entanglement source
Figure 5: Zoom-in view of the “Entanglement source” part of the experimental setup shown in Fig. 2a in main text for generating maximally entangled photon pairs.
@@
For our experiment, the power intensity of pump light is set at $5$ mW, and we observe a twofold coincidence count rate of 73 kHz. The two photons are coupled out by two collimators, and the polarization of each photon is corrected by polarization controller (PC) consisting two QWPs and a HWP. Finally, we let photon B go through an additional HWP at 0 to produce the singlet state $\varrho_{0}^{H,V}=(|\Psi^{-}\rangle\!\langle\Psi^{-}|)_{\text{AB}}$, where $|\Psi^{-}\rangle_{\text{AB}}=\frac{1}{\sqrt{2}}|H\rangle|V\rangle-|V\rangle|H\rangle)_{\text{AB}}$.
@@ A.5 Experimental realization of an arbitrary Werner state
Although the decomposition that we have provided in Eq. (2) in main text suffices for our experimental demonstration, it is not valid for $v>\frac{d+1}{2d}$, which leads to a negative value of $q=1-\frac{2d}{d+1}v$. For completeness, we provide below an alternative two-qudit decomposition of $W^{(d)}(v)$ valid for all $v\in[0,1]$. Explicitly, we note that
@@
A.6 ScLFs beyond qubit projections
Although we have focused on the projection onto a qubit subspace in our experimental demonstration, it is worth noting that other ScLFs are possible with the path encoding of quantum information that we have employed. For example, suppose we have encoded our two-qudit quantum state $\rho$ using locally the $d$ path states $\{|l_{k}\rangle\}_{k=0}^{d-1}$ and we want to perform a local projection onto the subspace $\mathcal{H}_{\mathcal{D}}=\text{span}\{|l_{l}\rangle\}_{k\in\mathcal{D}}$ where $\mathcal{D}\subset\bm{D}:=\{0,1,\cdots,d-1\}$, then it suffices to block the paths corresponding to $\bm{D}\setminus\mathcal{D}$. Then, if there are $d^{\prime}$ elements in $\mathcal{D}$, such a filtering operation will transform locally the state space from dimension $d$ to dimension $d^{\prime}$. In these notations, the specific ScLFs described in the main text simply correspond to $d=3$ and $\mathcal{D}=\{1,2\}$ with $d^{\prime}=2$.
More generally, a local filtering operation transforming a $d$-dimensional state to a $d^{\prime}$-dimensional state is described by a rectangular matrix $M$ of size $d^{\prime}\times d$ such that $\|M\|_{\infty}=1$ (i.e., having its maximum singular value equal to 1). Any such matrix $M$ admits the singular value decomposition
$M=U\Sigma V^{\dagger}$ (14)
where $V$ is a $d\times d$ unitary matrix, $\Sigma=\text{diag}(s_{0},s_{1},\cdots)$ is a $d^{\prime}\times d$ matrix consisting of the singular values $s_{i}$ of $M$ on its diagonal, and $U$ is a $d^{\prime}\times d^{\prime}$ unitary matrix. To implement the local filtering operation corresponding to $M$, it then suffices to perform the following steps in sequence:
  1. 1.
    Apply the unitary transformation $V^{\dagger}$ on the unfiltered space spanned by $\{|l_{k}\rangle\}_{k=0}^{d-1}$.
  2. 2.
    Apply the projection onto $\mathcal{H}_{\mathcal{D}}$ where $\mathcal{D}$ consists of the collection of indices in $\bm{D}$ such that $s_{i}\neq 0$ for all $i\in\mathcal{D}$. More precisely, in addition to blocking the paths corresponding to $\bm{D}\setminus\mathcal{D}$, signals propagating through path $i$ are also attenuated by the factor $s_{i}$.
  3. 3.
    Apply the unitary transformation $U$ on the filtered space spanned by $\{|l_{k}\rangle\}_{k\in\mathcal{D}}$.
Note that the experimental preparation of good quality high-dimensional states and their distillation are still challenging with bulk optics, especially for large $d$. However, it is worth noting that integrated optics offers a promising and scalable approach to generate $d$-dimensional entangled photons encoded in the path degree of freedom Wang et al. (2018); Bao et al. (2023). In fact, the general single-copy local filtering operations discussed above can be implemented on a single chip in this physical implementation. More precisely, arbitrary unitary operations $U$ and $V$ of Eq. 14 can be implemented with $d^{\prime}$-channel and $d$-channel Mach-Zehnder interferometer (MZI) network Reck et al. (1994); Clements et al. (2016). The blocking of path $i$, as well as the attenuation $s_{i}$ can also be implemented by a programmable MZI, where the output can be thermally or electrically tuned.
@@ Appendix B Miscellaneous Details and Theoretical Results
B.1 Fully-entangled fraction
In the main text, we claim the following Proposition
Proposition 1.
If a two-qubit state $\rho$ satisfies $F_{2}>\frac{1}{2}$, it must also satisfy $F_{d}>\frac{1}{d}$ for all integer $d>2$.
In the following, we give a proof of this Proposition.
Proof.
Let $\rho$ be a two-qubit state acting on the local qubit space $\mathcal{H}_{2}=$ span$\{|0\rangle,|1\rangle\}$ and $|\Psi^{\prime}_{2}\rangle$ be the maximally entangled state (MES) saturating the fully-entangled fraction of $\rho$, i.e.,
$F_{2}(\rho)=\langle\Psi^{\prime}_{2}|\rho|\Psi^{\prime}_{2}\rangle\quad\text{with}\quad|\Psi^{\prime}_{2}\rangle=\mathbb{I}_{2}\otimes U_{2}|\Phi^{+}_{2}\rangle$ (15)
and $|\Phi^{+}_{d}\rangle:=\frac{1}{\sqrt{d}}\sum_{i=0}^{d-1}|i\rangle|i\rangle$.
Now, define the $d\times d$ unitary operator $U_{d}:=U_{2}\oplus\mathbb{I}_{d-2}$, where the identity operator $\mathbb{I}_{d-2}$ acts on span$\{|2\rangle,|3\rangle,\cdots,|d-1\rangle\}$. Let $|\Psi^{\prime}_{d}\rangle:=\mathbb{I}_{d}\otimes U_{d}|\Phi^{+}_{d}\rangle$, then by embedding $\rho$ into the two-qudit space as $\tilde{\rho}=\rho\oplus{\bf 0}_{d^{2}-4}$, with the null operator ${\bf 0}_{d^{2}-4}$ acting on the subspace of $\mathbb{C}^{d}\otimes\mathbb{C}^{d}$ orthogonal to $\mathcal{H}_{2}\otimes\mathcal{H}_{2}$, we get
$F_{d}(\rho)$ $=F_{d}(\tilde{\rho}):=\max_{\text{MES\ }|\Psi_{d}\rangle}\langle\Psi_{d}|\tilde{\rho}|\Psi_{d}\rangle$
$\geq\langle\Psi^{\prime}_{d}|\tilde{\rho}|\Psi^{\prime}_{d}\rangle=\langle\Phi^{+}_{d}|(\mathbb{I}_{d}\otimes U_{d}^{\dagger})\tilde{\rho}(\mathbb{I}_{d}\otimes U_{d})|\Phi^{+}_{d}\rangle$
$=\langle\Phi^{+}_{d}|(\mathbb{I}_{d}\otimes U_{d}^{\dagger})(\rho\oplus{\bf 0}_{d^{2}-4})\times$
$\qquad\qquad\qquad\qquad\left[(\mathbb{I}_{2}\oplus\mathbb{I}_{d-2})\otimes(U_{2}\oplus\mathbb{I}_{d-2})|\Phi^{+}_{d}\rangle\right]$
$=\langle\Phi^{+}_{d}|(\mathbb{I}_{d}\otimes U_{d}^{\dagger})\left\{\left[\rho~{}(\mathbb{I}_{2}\otimes U_{2})\right]\oplus{\bf 0}_{d^{2}-4}\right\}|\Phi^{+}_{d}\rangle$
$=\langle\Phi^{+}_{d}|\left\{\left[(\mathbb{I}_{2}\otimes U_{2}^{\dagger})\,\rho\,(\mathbb{I}_{2}\otimes U_{2})\right]\oplus{\bf 0}_{d^{2}-4}\right\}|\Phi^{+}_{d}\rangle$
$=\frac{1}{d}\sum_{i,j=0}^{d-1}\langle i|\langle i|\!\left\{\left[(\mathbb{I}_{2}\otimes U_{2}^{\dagger})\,\rho\,(\mathbb{I}_{2}\otimes U_{2})\right]\!\oplus{\bf 0}_{d^{2}-4}\right\}\!|j\rangle|j\rangle$
$=\frac{1}{d}\sum_{i,j=0}^{1}\langle i|\langle i|\!\left[(\mathbb{I}_{2}\otimes U_{2}^{\dagger})\,\rho\,(\mathbb{I}_{2}\otimes U_{2})\right]\!|j\rangle|j\rangle$
$=\frac{2}{d}\langle\Phi^{+}_{2}|(\mathbb{I}_{2}\otimes U_{2}^{\dagger})\,\rho\,(\mathbb{I}_{2}\otimes U_{2})|\Phi^{+}_{2}\rangle$
$=\frac{2}{d}\langle\Psi_{2}^{\prime}|\rho|\Psi_{2}^{\prime}\rangle>\frac{2}{d}\times\frac{1}{2}=\frac{1}{d}$
where we have used in the last line Eq. 15 and the premise $F_{2}(\rho)>\frac{1}{2}$. ∎
B.2 Bell-nonlocality
B.2.1 Symmetric (quasi-)extension
@@
For completeness (and simplicity of presentation), we provide below a semidefinite program (SDP) that facilitates the search for the existence of a $(k,1)$-symmetric extension Doherty et al. (2002) for any given $\rho_{\text{AB}}$ acting on $\mathbb{C}^{d_{A}}\otimes\mathbb{C}^{d_{B}}$. For the rest of this subsection, we shall refer to any $\rho_{\text{AB}}$ admitting a $(k,1)$-symmetric extension as being $k$-extendible. Our SDP formulation differs somewhat from that of Doherty et al. (2002); Terhal et al. (2003) but makes explicit use of the two facts: (1) the set of $k$-extendible density matrices $\rho_{\text{AB}}$ is convex, and (2) the maximally mixed state $\frac{\mathbb{I}_{D}}{D}$, where $D:=d_{A}d_{B}$ is $k$-extendible for all $k>1$. Hence, if mixing $\rho_{\text{AB}}$ with a negatively-weighted $\frac{\mathbb{I}_{D}}{D}$ is $k$-extendible, $\rho_{\text{AB}}$ itself must already be $k$-extendible:
@@
For Werner states $W^{(d)}(w)$, using the identity $\Pi^{d}_{+}=\mathbb{I}-\Pi^{d}_{-}$, we can also rewrite it from Eq. (1) in main text as
$v_{t}=\frac{n^{d}_{+}}{D}w_{t}=\frac{n^{d}_{+}}{D}\frac{t^{\star}_{0}-1}{t^{\star}_{0}},$ (20)
i.e., $W^{(d)}(v)$ for all $v>v_{t}$ is $k$-extendible.
$\ \ \min_{t\geq 0}\qquad\qquad\qquad\qquad t$
@@
From Table 2, we see that the tomographically-reconstructed experimentally states $\rho_{\text{expt}}(v)$ more or less exhibit the same level of $k$-(quasi)extendability for $k\leq 4$. For larger values of $k$, we remark that by using an adaptation Zhang of the method from Villegas-Aguilar et al. (2024), we could verify that for $v=0.45$, a convex decomposition of the following kind can be obtained for the experimentally produced state:
$\begin{split}\rho_{\text{expt}}(v)&=p\,W^{(3)}(v)+\sum_{i}q_{i}\,\mathbb{I}\otimes U_{i}W^{(3)}(v)\mathbb{I}\otimes U_{i}^{\dagger}\\ &+\sum_{\ell,j,m}r_{\ell jm}\,V_{\ell}\otimes U_{\ell}|j\rangle\!\langle j|\otimes|m\rangle\!\langle m|\,V_{\ell}^{\dagger}\otimes U_{\ell}^{\dagger},\end{split}$ (22)
where $\{|j\rangle\}$ and $\{|m\rangle\}$ are computational basis states for the first and second Hilbert spaces, $V_{\ell}$ and $U_{\ell}$ are some unitaries acting on the two Hilbert space, $p,q_{i},r_{\ell jm}\geq 0$ for all $i,\ell,j,m$, and $p+\sum_{i}p_{i}+\sum_{\ell jk}r_{\ell jm}=1$. Notice that each term in the first sum on the RHS of Eq. 22 has exactly the same (non)local property as $W^{(3)}(v)$ while the second sum clearly gives a separable state, which is $k$-(quasi)extendable for all integers $k\geq 1$. Using the convexity of the set of $k$-(quasi)extendable states (for any fixed but arbitrary $k$), we thus see that if $W^{(3)}(v)$ is $k$-(quasi)extendable for some $k$, so is the corresponding $\rho_{\text{expt}}(v)$.
Note further that unlike the case of two-qubit states discussed in Villegas-Aguilar et al. (2024), the partial-transposition criterion Horodecki et al. (2009) is only necessary but not sufficient for separability for the higher-dimensional systems we are considering. Consequently, in searching for a decomposition like Eq. 22 for a given $v$, we settle for a restricted set of separable states given by the second sum of Eq. 22 (facilitated by a randomly sampled set of unitaries), rather than optimizing over all separable two-qutrit states.
B.2.2 Finding Bell-nonlocal correlations for the experimentally produced two-qutrit states